# A Game-Theoretic Unital Classification Theorem for $C^{*}$-Algebras

### v1

@@ 2.1. Descriptive set theory preliminaries
We give here a brief summary of some standard definitions and facts from descriptive set theory that will be relevant throughout. For a more thorough treatment of the subject see [MR1321597, MR561709].
@@
A subset of a standard Borel space is called analytic if it is the image of a Borel subset by a Borel map between standard Borel spaces. Equivalently, analytic subsets can be defined as projections of closed (or Borel) subsets of a product with another standard Borel space. A fact that we will use throughout is that analytic sets are closed under countable unions and intersections, and images and preimages of analytic sets under Borel functions are analytic. For $X$ a standard Borel space and $X_{0}\subseteq X$, we say that a set $A\subseteq X_{0}$ is analytic (resp. Borel) relative to $X_{0}$ if it is the intersection of an analytic (resp. Borel) subset of $X$ with $A$. Since analytic (resp. Borel) sets are closed under finite intersections, if $X_{0}$ is analytic then a set is analytic relative to $X_{0}$ if and only if it is analytic in $X$.
@@
Consequently, a map between standard Borel spaces whose graph is analytic (as a subset of the product) is a Borel function (with a posteriori Borel graph).
@@ 2.2. Infinitary metric logic and the space of separable structures
Given a countable, discrete relational language $L=\{R_{i}~\colon i\in I\}$, each countably infinite $L$-structure is isomorphic to an $L$-structure $A=(\omega;(R_{i}^{A})_{i\in I})$ on $\omega$. This structure $A$ can be viewed as an element of the set $\operatorname{Mod}_{\omega}(L):=\prod_{i\in I}2^{\omega^{n_{i}}}$ where $n_{i}$ denotes the arity of $R_{i}$. As a countable product of finite powers of the Cantor space, $\operatorname{Mod}_{\omega}(L)$ carries a natural Polish topology, which coincides with the topology generated by open sets of the form $\operatorname{Mod}_{\omega}(\varphi(a)):=\{A\colon A\models\varphi(a)\}$ where $\varphi$ is a quantifier-free formula. This leads to a fruitful interplay between (infinitary) first-order logic and descriptive set theory that has been studied extensively. Sets of the form $\operatorname{Mod}_{\omega}(\sigma)$ where $\sigma$ is an $L_{\omega_{1},\omega}$-formula are Borel in $\operatorname{Mod}_{\omega}(L)$, as can be seen by induction. A theorem of Lopez-Escobar [Lopez1965] establishes that conversely, if $X\subseteq\operatorname{Mod}_{\omega}(L)$ is Borel and invariant under the action of $S_{\infty}$, then $X=\operatorname{Mod}_{\omega}(\sigma)$ for some $L_{\omega_{1},\omega}$-sentence $\sigma$ (see also [MR1321597, Chapter II.16]). In a similar direction, Scott’s isomorphism theorem for countable structures produces for each $A\in\operatorname{Mod}_{\omega}(L)$ an $L_{\omega_{1},\omega}$-sentence $\sigma_{A}$ called the Scott sentence of $A$ such that the isomorphism class of $A$ in $\operatorname{Mod}_{\omega}(L)$ coincides with $\operatorname{Mod}(\sigma_{A})$.
@@
In view of our main application in Section 3, we are interested in the more general case of a continuous logic language $L$, and the study of spaces of separable $L$-structures. Roughly speaking, continuous logic extends discrete first-order logic by allowing relations – and therefore all formulas – to take real values and allowing as connectives arbitrary continuous functions. For a detailed treatment of continuous logic, we refer the reader to [MR2436146, MR4654490].
@@
A key technical subtlety that arises in moving to such a logic is that in order to obtain a compactness theorem, one needs to ensure that each formula takes values only in some bounded interval, uniformly in each structure. We use here the framework of metric logic, as developed in [MR2436146], where every language $L$ contains a distinguished binary predicate symbol $d$ to be interpreted as a metric, and an $L$-structure is a metric space, with each relation symbol $R$ taking values in a compact interval $I_{R}\subseteq\mathbb{R}$, determined by $L$. In particular, the language $L$ requires each structure to have diameter bounded by some fixed $D_{L}>0$. It is natural to insist functions and relations be continuous on a given structure with respect to the metric. To ensure that continuity passes through ultraproducts, one demands each function and relation respect some moduli of uniform continuity. Given $\Delta\colon[0,\infty)^{n}\to[0,\infty)$ and a metric space $X$, we define for $x,x^{\prime}\in X^{n}$,
@@
Let $L$ be a metric language and assume for notational simplicity that $L$ is one-sorted. An $L$-structure A{A} consists of a complete metric space $(A,d_{A})$ with diameter $\leq D_{L}$, together with
@@
L{L}-terms are defined inductively from constant and function symbols as in discrete logic, and each $L$-term $\tau$ carries a modulus of continuity $\Delta_{\tau}$, obtained inductively from those on function symbols. Atomic formulas are of the form $R(\tau_{1},\dots,\tau_{n})$ where $\tau_{i}$ are terms and $R$ is a relation symbol. L{L}-formulas are built inductively from atomic formulas, with arbitrary uniformly continuous maps $u\colon\mathbb{R}^{n}\to\mathbb{R}$ as connectives, and quantifiers $\inf_{x}$ and $\sup_{x}$ (in place of $\exists x$ and $\forall x$). Formulas are interpreted in L{L}-structures in the canonical way.
We move now to infinitary metric logic following [Ben_Yaacov_2009]. A technical hurdle of considering infinitary formulas in metric languages is that when taking infima and suprema (in place of the discrete infinitary connectives $\bigvee$ and $\bigwedge$) over countable families of formulas, the resulting formula may fail to respect any modulus of continuity or have unbounded range. This motivates the restriction on infinitary conjunctions and disjunctions in the fourth point of the following definition.
Fix a metric language $L$. Note that the intervals $I_{R}$ and moduli of continuity $\Delta_{R}$ respectively $\Delta_{\tau}$ on relation symbols respectively terms in $L$ can be extended inductively in a canonical way to compact intervals $I_{\varphi}$ and moduli of uniform continuity $\Delta_{\varphi}$, when $\varphi$ is an infinitary formula as in the below definition.
@@
    If $\varphi\in L_{\omega_{1},\omega}$ and $x$ is a variable then $\sup_{x}\varphi,~\inf_{x}\varphi\in L_{\omega_{1},\omega}$ with
@@
    Given variables $x_{1},\dots,x_{n}$, a bounded interval $I\subseteq\mathbb{R}$ and a modulus of continuity $\Delta\colon[0,\infty)^{n}\to[0,\infty)$, if $(\varphi_{i})_{i<\omega}$ is such that for each $i<\omega$, $\varphi_{i}$ is an $L_{\omega_{1},\omega}$-formula with free variables $x_{1},\dots,x_{n}$, taking values in $I$ and respecting $\Delta$ (i.e., $\Delta_{\varphi_{i}}\leq\Delta$), then $\bigwedge_{i}\varphi_{i},\bigvee_{i}\varphi_{i}\in L_{\omega_{1},\omega}$ with
@@
Note that by definition, $\operatorname{qr}(\varphi)$ is a countable ordinal for any $\varphi\in L_{\omega_{1},\omega}$. The first-order $L$-formulas are the $L_{\omega_{1},\omega}$-formulas that do not use (iv) in their inductive definition. A sentence is a formula without free variables. A (first-order) theory $T$ is a set $T=\{\varphi_{i}\colon i\in I\}$ for some (first-order) sentences $\{\varphi_{i}\}_{i\in I}$. For an $L_{\omega_{1},\omega}$-sentence $\sigma$ and an $L$-structure AA, we write $\sigma^{A}$ for the value of $\sigma$ interpreted in $A$. We say that $A$ satisfies $\sigma$ (written $A\models\sigma$), if $\sigma^{A}=0$. Similarly, for a theory $T$, we write $A\models T$ if $A$ satisfies all sentences from $T$. In that case we call $A$ a model of $T$.
We define the language and theory of $C^{*}$-algebras below for use in future applications in Section 3.
@@
Consider the multi-sorted countable language $L_{C^{*}}$ for $C^{*}$-algebras, with:
  • for each $n\in\mathbb{N}$, sorts $B_{n}$ to be interpreted as the ball of radius $n$, predicates $d_{n}\colon B_{n}\to[0,n]$,
@@
    • $_{n}\colon B_{n}\times B_{n}\to B_{2n}$,
    • ${{}^{*}}_{n}\colon B_{n}\to B_{n}$,
    • $\cdot_{n}\colon B_{n}\times B_{n}\to B_{n^{2}}$,
    • $(\cdot\lambda)_{n}\colon B_{n}\to B_{m}$ for $m={\lceil|\lambda|n\rceil}$ and $\lambda\in\mathbb{Q}(i)$,
    • $I_{mn}\colon B_{m}\to B_{n}$ for $m<n$,
  • and a constant symbol $0\in B_{1}$.
Furthermore, we write $T_{C^{*}}$ for the theory of $C^{*}$-algebras as presented in Example 2.2.1 from [farah2021MTofC*], except we only use $\mathbb{Q}(i)$ scalars (see loc.cit. for a natural choice of moduli of continuity and some discussion). We also denote by $T_{C^{*}}^{\prime}$ the theory of unital $C^{*}$-algebras in the language $L_{C^{*}}$ expanded by a constant symbol $1\in B_{1}$.
@@
Consider the multi-sorted language $L_{ous}$ (“ousous” stands for order unit spaces) consisting of:
  • for each $n\in\mathbb{N}$, sorts $B_{n}$ with metric predicates $d_{n}\colon B_{n}\to[0,n]$,
@@
Definition 2.6.
We will be interested throughout this paper in separable metric structures. Let ${L}=(d;(R_{i})_{i};(f_{j})_{j};c_{k})$ be a one-sorted countable metric language. Since $L$ is countable, any separable $L$-structure AA contains a countable dense subset $A_{0}\subseteq A$, containing all $c_{k}^{A}$ and closed under all $f_{j}^{A}$. Such $A_{0}$ is sometimes called a pre-substructure, and a substructure if additionally, $A_{0}$ is closed as a metric space. In this case, $A$ is uniquely determined by $A_{0}$.
@@
with the standard Borel structure induced by the product topology of the respective Polish topologies on $\mathbb{R}^{\omega^{n}}$ and $\omega$. Then we can assign to $A$ an element $A\in X_{L}$ (by abuse of notation) in the canonical way, setting
@@
We define $\operatorname{Mod}_{\omega}(L)\subseteq X_{L}$ to be the range of this assignment. We assume in the definition below that $L$ is a countable one-sorted metric language as in Definition 2.2. Note that the construction can be generalized straightforwardly to a language with countably many sorts. For example, a separable $L_{C^{*}}$-structure $A$ can be encoded as an element of the space
(2) $\prod_{n<\omega}\Big(\mathbb{R}^{\omega^{2}}\times\omega^{\omega^{2}}\times\omega^{\omega^{2}}\times\omega^{\omega}\times\omega^{\omega\times\mathbb{Q}(i)}\Big)\times\big(\prod_{m<n<\omega}\omega^{\omega}\big)\times\omega^{\omega\times\omega}\times\omega\times\omega$
by fixing dense countable subsets of each ball $B_{n}$ and closing them under all functions. Recall that $L$ specifies a diameter $D>0$.
@@
By the metric generalization of Scott’s isomorphism theorem proved in [yaacov2017metric, Theorem 1.2], the isomorphism class in $\operatorname{Mod}_{\omega}(L)$ of any separable $L$-structure AA is of the form $\operatorname{Mod}_{\omega}(\sigma_{A})$ for some $\sigma_{A}\in L_{\omega_{1},\omega}$. As a corollary, one obtains the following.
Fact 2.9 ([yaacov2017metric, Theorem 1.2]).
@@ 2.3. Metric dynamic Ehrenfeucht-Fraïssé games
An illustrative way of characterizing the expressive power of an (infinitary) logic are the so-called Ehrenfeucht-Fraïssé games. Going back to Hausdorff and what is now a standard back-and-forth argument for Cantor’s isomorphism theorem [hausdorff_grundzuege, Section 4.7], these games were first considered explicitly by Ehrenfeucht [ehrenfeucht1957application], based on the work of Fraïssé [fraisse1955quelques].
@@
Denoting by $\operatorname{EF}_{\omega}(A,B)$ the analogous game with $\omega$-many rounds, II has a winning strategy in $\operatorname{EF}_{\omega}(A,B)$ if and only if the structures are indistinguishable to the infinitary logic $L_{\omega_{1},\omega}$ [^5] In fact, this game corresponds to $L_{\infty\omega}$, but we are interested only in countable languages where these two logics coincide., i.e., $A\equiv_{\omega_{1}}B$. To capture the finer notion $\equiv_{\alpha}$ of elementary equivalence up to formulas of quantifier-rank $\alpha$ for some $\alpha<\omega_{1}$, one considers a version of the game with dynamic length. The dynamic Ehrenfeucht-Fraïssé game $\operatorname{EFD}_{\alpha}(A,B)$ is played like $\operatorname{EF}_{n}(A,B)$, with the following modification. There is a ‘game-clock’, call it $\beta$, that is set to $\alpha$ before the game starts. In each round, Player I, in addition to playing a usual move, sets the game-clock to some ordinal $\beta^{\prime}<\beta$. The game ends when $\beta=0$. Note that if $\alpha\geq\omega$, a play of the game can be arbitrarily long, but will always be finite. The advantage granted to Player I in the game clock amounts exactly to the expressive power gained in increasing quantifier rank: II has a winning strategy in $\operatorname{EFD}_{\alpha}(A,B)$ if and only if $A\equiv_{\alpha}B$. This is due to Karp [KARP2014407].
An analogue of $L_{\omega_{1}\omega}$ for metric languages $L$ was developed in [yaacov2017metric], where a version of both the Lopez-Escobar and the Scott Theorem is proved, and a metric Scott analysis is developed. We give the setup and main theorems of [yaacov2017metric] here.
@@
We define in this section a generalization of the dynamic Ehrenfeucht-Fraïssé game for metric languages based on the back-and-forth pseudo-distance of [yaacov2017metric], and show in Theorem 2.21 that the resulting refining sequence of equivalence relations $(\equiv^{bf}_{\alpha})_{\alpha<\omega_{1}}$ is uniformly Borel in the sense of Defintion 2.25. The naïve approach to defining metric EF-games would be to play exactly as in the discrete case, with II winning if $\sup_{\varphi}|\varphi^{A}(a)-\varphi^{B}(b)|=0$ where $\varphi$ ranges over all finitary, quantifier-free formulas. This game is too strong, for several reasons. For one, when we consider separable structures $A$ and $B$, we want to restrict the legal moves to some previously fixed dense countable substructures of $A$ and $B$, without changing the outcome of the game. The above described game is sensitive to this restriction: Take $A\in\operatorname{Mod}_{\omega}(L_{C^{*}})$ a separable $C^{*}$-algebra, with a countable dense pre-substructure $A_{0}\subseteq A$. Find $r\in\mathbb{R}\setminus\mathbb{Q}$ such that $a||\neq r$ for all $a\in A_{0}$. Setting $A_{0}^{\prime}$ to be the (still countable) pre-substructure generated by $A_{0}\cup\{r\}$, we see that Player I can win in one round in a game between $A_{0}$ and $A_{0}^{\prime}$ – both of which have completion $A$ – by playing $r\in A_{0}^{\prime}$. We would like to remedy this by loosening the winning requirement for Player II such that values of quantifier-free formulas may differ up to some $\varepsilon>0$ on the played tuples. Note however that this only makes sense if we impose some restriction on the formulas we consider. If $a\in A^{n}$, $b\in B^{n}$ are such that $|\varphi^{A}(a)-\varphi^{B}(b)|>0$ for some quantifier-free $\varphi$, then
$\sup_{\psi}|\psi^{A}(a)-\psi^{B}(b)|\geq\sup_{n}|(n\varphi)^{A}(a)-(n\varphi)^{B}(b)|=\infty$
@@
where $\psi$ ranges over all quantifier-free formulas. Following [yaacov2017metric], this is dealt with by restricting only to formulas $\psi$ which respect a fixed, so-called weak modulus $\Omega\colon[0,\infty)^{\mathbb{N}}\to[0,\infty]$ in the sense of [yaacov2017metric, Definition 2.2]. That is, the truncations
@@
$\Omega|_{n}\colon\begin{cases}[0,\infty)^{n}\to[0,\infty)\\ (\delta_{1},\dots,\delta_{n})\mapsto(\delta_{1},\dots,\delta_{n},0,0\dots)\end{cases}$
@@
Definition 2.10 ([yaacov2017metric, Definition 3.1]).

### v2

@@ 2.1. Descriptive set theory preliminaries
We give here a brief summary of some standard definitions and facts from descriptive set theory that will be relevant throughout. For standard references on the subject, see [39, 45].
@@
A subset of a standard Borel space is called analytic if it is the image of a Borel subset by a Borel map. Equivalently, analytic subsets can be defined as projections of closed (or Borel) subsets of a product with another standard Borel space. A fact that we will use throughout is that analytic sets are closed under countable unions and intersections, and images and preimages of analytic sets under Borel functions are analytic. For a standard Borel space $X$ and $X_{0}\subseteq X$, we say that a set $A\subseteq X_{0}$ is analytic (resp. Borel) relative to $X_{0}$ if it is the intersection of an analytic (resp. Borel) subset of $X$ with $A$. Since analytic sets are closed under finite intersections, if $X_{0}$ is analytic then a set is analytic relative to $X_{0}$ if and only if it is analytic in $X$.
@@
As a consequence of Suslin’s theorem, a map between standard Borel spaces whose graph is analytic (as a subset of the product) is a Borel function (with a posteriori Borel graph).
@@ 2.2. Infinitary metric logic and the space of separable structures
Given a countable, discrete relational language $L=\{R_{i}~\colon i\in I\}$, each countably infinite $L$-structure is isomorphic to an $L$-structure $A=(\omega;(R_{i}^{A})_{i\in I})$ on $\omega$. This structure $A$ can be viewed as an element of the set $\operatorname{Mod}_{\omega}(L):=\prod_{i\in I}2^{\omega^{n_{i}}}$ where $n_{i}$ denotes the arity of $R_{i}$. As a countable product of finite powers of the Cantor space, $\operatorname{Mod}_{\omega}(L)$ carries a natural Polish topology, which coincides with the topology generated by open sets of the form $\operatorname{Mod}_{\omega}(\varphi(a)):=\{A\colon A\models\varphi(a)\}$ where $\varphi$ is a quantifier-free formula. This leads to a fruitful interplay between (infinitary) first-order logic and descriptive set theory that has been studied extensively. Sets of the form $\operatorname{Mod}_{\omega}(\sigma)$ where $\sigma$ is an $L_{\omega_{1},\omega}$-formula are Borel in $\operatorname{Mod}_{\omega}(L)$, as can be seen by induction. A theorem of Lopez-Escobar [43] establishes that conversely, if $X\subseteq\operatorname{Mod}_{\omega}(L)$ is Borel and invariant under the action of $S_{\infty}$, then $X=\operatorname{Mod}_{\omega}(\sigma)$ for some $L_{\omega_{1},\omega}$-sentence $\sigma$ (see also [39, Chapter II.16]). Scott’s isomorphism theorem for countable structures produces for each $A\in\operatorname{Mod}_{\omega}(L)$ an $L_{\omega_{1},\omega}$-sentence $\sigma_{A}$ called the Scott sentence of $A$ such that the isomorphism class of $A$ in $\operatorname{Mod}_{\omega}(L)$ coincides with $\operatorname{Mod}(\sigma_{A})$.
@@
In view of our main application in Section 3, we are interested in the more general case of a continuous logic language $L$, and the study of spaces of separable $L$-structures. Roughly speaking, continuous logic extends discrete first-order logic by allowing relations – and therefore all formulas – to take real values and allowing as connectives arbitrary continuous functions. For a detailed treatment of continuous logic, we refer the reader to [1, 29].
@@
A key technical subtlety that arises in moving to such a logic is that in order to obtain a compactness theorem, one needs to ensure that each formula takes values only in some bounded interval, uniformly in each structure. We use here the framework of metric logic, as developed in [1], where every language $L$ contains a distinguished binary predicate symbol $d$ to be interpreted as a metric, and an $L$-structure is a complete metric space, with each relation symbol $R$ taking values in a compact interval $I_{R}\subseteq\mathbb{R}$, determined by $L$. In particular, the language $L$ requires each structure to have diameter bounded by some fixed $D_{L}>0$. It is natural to insist functions and relations be continuous on a given $L$-structure with respect to the metric. To ensure that continuity passes through ultraproducts, one demands each function and relation respect some modulus of uniform continuity. Given $\Delta\colon[0,\infty)^{n}\to[0,\infty)$ and a metric space $X$, we define for $x,x^{\prime}\in X^{n}$,
@@
Let $L$ be a metric language and assume for notational simplicity that $L$ is one-sorted, that is $|S|=1$. An $L$-structure A{A} consists of a complete metric space $(A,d_{A})$ with diameter $\leq D_{L}$, together with
@@
L{L}-terms are defined inductively from constant and function symbols as in discrete logic, and each $L$-term $\tau$ carries a modulus of continuity $\Delta_{\tau}$, obtained inductively from those on function symbols. Atomic formulas are of the form $R(\tau_{1},\dots,\tau_{n})$ where $\tau_{i}$ are terms and $R$ is a relation symbol. L{L}-formulas are built inductively from atomic formulas, with arbitrary uniformly continuous maps $u\colon\mathbb{R}^{n}\to\mathbb{R}$ as connectives, and quantifiers $\inf_{x}$ and $\sup_{x}$ (in place of $\exists x$ and $\forall x$).
We move now to infinitary metric logic following [3]. A technical hurdle of considering infinitary formulas in metric languages is that when taking infima and suprema (in place of the discrete infinitary connectives $\bigvee$ and $\bigwedge$) over countable families of formulas, the resulting formula may fail to be continuous or have unbounded range. This motivates the restriction on infinitary conjunctions and disjunctions in the fourth point of Definition 2.3 below.
Fix a metric language $L$. Note that the intervals $I_{R}$ and moduli of continuity $\Delta_{R}$ respectively $\Delta_{\tau}$ on relation symbols respectively terms in $L$ can be extended inductively in a canonical way to compact intervals $I_{\varphi}$ and moduli of uniform continuity $\Delta_{\varphi}$, when $\varphi$ is an infinitary formula as in the below definition. We write $\varphi\in L_{\omega_{1},\omega}$ when $\varphi$ is an $L_{\omega_{1},\omega}$-formula. Recall that $\omega_{1}$ denotes the smallest uncountable ordinal.
@@
    If $\varphi\in L_{\omega_{1},\omega}$ and $x$ is a variable then $\sup_{x}\varphi\in L_{\omega_{1},\omega}$ and $\inf_{x}\varphi\in L_{\omega_{1},\omega}$ with
@@
    Given variables $x_{1},\dots,x_{n}$, a bounded interval $I\subseteq\mathbb{R}$ and a modulus of continuity $\Delta\colon[0,\infty)^{n}\to[0,\infty)$, if $(\varphi_{i})_{i<\omega}$ is such that for each $i<\omega$, $\varphi_{i}$ is an $L_{\omega_{1},\omega}$-formula with free variables $x_{1},\dots,x_{n}$, taking values in $I$ and respecting $\Delta$ (i.e., $\Delta_{\varphi_{i}}\leq\Delta$), then $\bigwedge_{i}\varphi_{i}\in L_{\omega_{1},\omega}$ and $\bigvee_{i}\varphi_{i}\in L_{\omega_{1},\omega}$ with
@@
Note that by definition, $\operatorname{qr}(\varphi)$ is a countable ordinal (written $\operatorname{qr}(\varphi)<\omega_{1}$) for any $\varphi\in L_{\omega_{1},\omega}$. The first-order $L$-formulas are those defined only by (i)(iii)(i)-(iii) in the above definition. A sentence is a formula without free variables. A (first-order) theory $T$ is a set $T=\{\varphi_{i}\colon i\in I\}$ for some (first-order) sentences $\{\varphi_{i}\}_{i\in I}$. For an $L_{\omega_{1},\omega}$-sentence $\sigma$ and an $L$-structure AA, we write $\sigma^{A}$ for the value of $\sigma$ interpreted in $A$. We say that $A$ satisfies $\sigma$ (written $A\models\sigma$), if $\sigma^{A}=0$. Similarly, for a theory $T$, we write $A\models T$ if $A$ satisfies all sentences from $T$. In that case we call $A$ a model of $T$. For $L$-structures AA and $B$, we write $A\cong B$ and say that $A$ and $B$ are isomorphic if there exists a bijection $f:A\to B$ with $\varphi^{A}(a)=\varphi^{B}(f(a))$ for each $L$-formula $\varphi(x)$ and tuple $a$ in $A$ of appropriate arity. The following gives a family of weaker equivalence relations between $L$-structures.
Definition 2.4.
We define the language and theory of $C^{*}$-algebras and of order unit spaces below for use in future applications in Section 3.
@@
Consider the multi-sorted countable language $L_{C^{*}}$ for $C^{*}$-algebras, containing for each $n\in\mathbb{N}$,
  • a sort $B_{n}$ to be interpreted as the ball of radius $n$ and a predicate $d_{n}\colon B_{n}\to[0,n]$ to be interpreted as the metric,
@@
    • $_{n}\colon B_{n}\times B_{n}\to B_{2n}$ to be interpreted as addition,
    • ${{}^{*}}_{n}\colon B_{n}\to B_{n}$ to be interpreted as involution,
    • $\cdot_{n}\colon B_{n}\times B_{n}\to B_{n^{2}}$ to be interpreted as multiplication,
    • $(\cdot\lambda)_{n}\colon B_{n}\to B_{m}$ for $\lambda\in\mathbb{Q}(i)$ and $m={\lceil|\lambda|n\rceil}$, to be interpreted as scalar multiplication,
    • $I_{mn}\colon B_{m}\to B_{n}$ for $m<n$, to be interpreted as containment,
  • a constant symbol 0 of sort $B_{1}$, to be interpreted as zero.
Furthermore, we write $T_{C^{*}}$ for the theory of $C^{*}$-algebras as presented in Example 2.2.1 from [14], except we only use $\mathbb{Q}(i)$ scalars to ensure $T_{C^{*}}$ is countable (see loc.cit. for a natural choice of moduli of continuity and some discussion). We denote by $T_{C^{*}}^{\prime}$ the theory of unital $C^{*}$-algebras in the language $L_{C^{*}}$ expanded by a constant symbol $1\in B_{1}$.
@@
Consider the multi-sorted language $L_{ous}$ (“ousous” stands for order unit spaces) containing for each $n\in\mathbb{N}$,
  • a sort $B_{n}$ with metric predicates $d_{n}\colon B_{n}\to[0,n]$,
@@
We will be interested throughout this paper in separable metric structures. Let ${L}=(d;(R_{i})_{i};(f_{j})_{j};c_{k})$ be a one-sorted countable metric language. Since $L$ is countable, any separable $L$-structure AA contains a countable dense subset $A_{0}\subseteq A$, containing all constants $c_{k}^{A}$ and closed under all functions $f_{j}^{A}$. Such $A_{0}$ is sometimes called a pre-substructure, and a substructure if additionally, $A_{0}$ is closed as a metric space. When $A_{0}$ is a dense pre-substructure, $A$ is uniquely determined by $A_{0}$.
@@
with the standard Borel structure induced by the product topology of the respective Polish topologies on $\mathbb{R}^{\omega^{n_{i}}}$, $\omega^{\omega^{m_{j}}}$ and $\omega$. Then we can assign to $A$ an element $A\in X_{L}$ (by abuse of notation) in the canonical way, setting
@@
We define $\operatorname{Mod}_{\omega}(L)\subseteq X_{L}$ to be the range of this assignment. We assume in the definition below that $L$ is a countable one-sorted metric language as in Definition 2.2. Note that the construction can be generalized straightforwardly to a language with countably many sorts.
@@
By the metric generalization of Scott’s isomorphism theorem proved in [2, Theorem 1.2], the isomorphism class in $\operatorname{Mod}_{\omega}(L)$ of any separable $L$-structure AA is of the form $\operatorname{Mod}_{\omega}(\sigma_{A})$ for some $\sigma_{A}\in L_{\omega_{1},\omega}$. As a corollary, one obtains the following.
Fact 2.9 ([2, Theorem 1.2]).
@@ 2.3. Metric dynamic Ehrenfeucht-Fraïssé games
An illustrative way of characterizing the expressive power of an (infinitary) logic comes from so-called Ehrenfeucht-Fraïssé games. Going back to Hausdorff and what is now a standard back-and-forth argument for Cantor’s isomorphism theorem [30, Section 4.7], these games were first considered explicitly by Ehrenfeucht [10], based on the work of Fraïssé [18].
@@
Denoting by $\operatorname{EF}_{\omega}(A,B)$ the analogous game with $\omega$-many rounds, II has a winning strategy in $\operatorname{EF}_{\omega}(A,B)$ if and only if the structures are indistinguishable to the infinitary logic $L_{\omega_{1},\omega}$ [^5] In fact, this game corresponds to $L_{\infty,\omega}$, but we are interested only in countable languages where $L_{\infty,\omega}$ and $L_{\omega_{1},\omega}$ always coincide., i.e., $A\equiv_{\omega_{1}}B$. To capture the finer notion $\equiv_{\alpha}$ of elementary equivalence up to formulas of quantifier-rank $\alpha$ for some $\alpha<\omega_{1}$, one considers a version of the game with dynamic length. The dynamic Ehrenfeucht-Fraïssé game $\operatorname{EFD}_{\alpha}(A,B)$ is played like $\operatorname{EF}_{n}(A,B)$, with the following modification. There is a ‘game-clock’, call it $\beta$, that is set to $\alpha$ before the game starts. In each round, Player I, in addition to playing a usual move, sets the game-clock to some ordinal $\beta^{\prime}<\beta$. The game ends when $\beta=0$. Note that if $\alpha\geq\omega$, a play of the game can be arbitrarily long, but will always be finite since $\alpha$ is well-ordered. The advantage granted to Player I in the game clock amounts exactly to the expressive power gained in increasing quantifier rank: II has a winning strategy in $\operatorname{EFD}_{\alpha}(A,B)$ if and only if $A\equiv_{\alpha}B$. This is due to Karp [37].
@@
We define in this section a generalization of the dynamic Ehrenfeucht-Fraïssé game for metric languages based on the back-and-forth pseudo-distance of Ben Yaacov, Doucha, Nies and Tsankov [2], and show in Theorem 2.21 that the resulting refining sequence of equivalence relations $(\equiv^{bf}_{\alpha})_{\alpha<\omega_{1}}$ is uniformly Borel in the sense of Defintion 2.25. The naïve approach to defining metric EF-games would be to play exactly as in the discrete case, with II winning if $\sup_{\varphi}|\varphi^{A}(a)-\varphi^{B}(b)|=0$ where $\varphi$ ranges over all finitary, quantifier-free formulas. This game is too strong, however. When we consider separable structures $A$ and $B$, we want to restrict the legal moves to some previously fixed dense countable substructures of $A$ and $B$, without changing the outcome of the game. The above described game is sensitive to this restriction: Take $A\in\operatorname{Mod}_{\omega}(L_{C^{*}})$ to be a separable $C^{*}$-algebra, with a countable dense pre-substructure $A_{0}\subseteq A$. Find $r\in\mathbb{R}\setminus\mathbb{Q}$ such that $a||\neq r$ for all $a\in A_{0}$. Setting $A_{0}^{\prime}$ to be the (still countable) pre-substructure generated by $A_{0}\cup\{r\}$, we see that Player I can win in one round in a game between $A_{0}$ and $A_{0}^{\prime}$ – both of which have completion $A$ – by playing $r\in A_{0}^{\prime}$. We remedy this by loosening the winning requirement for Player II such that values of quantifier-free formulas may differ up to some $\varepsilon>0$ on the played tuples. Note however that this only makes sense if we impose some restriction on the formulas we consider. Indeed, if $a\in A^{n}$, $b\in B^{n}$ are such that $|\varphi^{A}(a)-\varphi^{B}(b)|>0$ for some quantifier-free $\varphi$, then
$\sup_{\psi}|\psi^{A}(a)-\psi^{B}(b)|\geq\sup_{n\in\mathbb{N}}|(n\varphi)^{A}(a)-(n\varphi)^{B}(b)|=\infty$
@@
where $\psi$ ranges over all quantifier-free formulas. Following [2], this is dealt with by restricting only to formulas $\psi$ which respect a fixed, so-called weak modulus $\Omega\colon[0,\infty)^{\mathbb{N}}\to[0,\infty]$ in the sense of [2, Definition 2.2]. That is, the truncations
@@
$\Omega|_{n}\colon\begin{cases}[0,\infty)^{n}\to[0,\infty)\\ (\delta_{1},\dots,\delta_{n})\mapsto\Omega(\delta_{1},\dots,\delta_{n},0,0\dots)\end{cases}$
@@
Definition 2.10 ([2, Definition 3.1]).